• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name flay

    Log entry time 07:55:11 on November 11, 2009

    Entry number 298957

    keyword=11/11/09 Wednesday Owl Shift Summary

    00:00 Start of shift running smoothly taking production data at I ~ 100 uA

    04:00 Production data taking going smoothly.

    05:45 Beam trip causes target temperature and pressure to drop. Same problem as last owl shift (11/10/09). Coda runs stopped. See HALOG 298938

    07:00 Production data taking running smoothly.

    07:27 Beam trip again causes target temp. & P to drop. We've decided this time to let the coda runs continue, as the target recovered quickly.

    07:50 Last run set started: 13881/25385/4938. Production running smoothly.

    List of Runs (PVDIS/LHRS/RHRS):


    13872/25375/4927 -- Stopped early because PVDIS DAQ froze; rebooted. Junk runs. 13873/25376/4928 13874/25377/4929 -- Stopped early. Kai changed the trigger prescales. See HALOG 298899. 13875/25378/4930 /25379/4931 /4932 13876/25380/4933 13877/25381/4934 13878/25382/4935 -- Stopped early. Beam trip caused the target pressure to drop (caused by a temperature drop) -- same issue as seen during last owl shift (11/10/09). See HALOG 298938. 13879/25383/4936 13880/25384/4937 -- Beam trip causes target temp & P to drop. We decided to leave the Coda running, as the trip was short and the target recovered quickly. 13881/25385/4938